OpenStudy (anonymous):

6x + 3x - 50 + x = 180 x=? A)x=10 B)x=23 C)x=20

OpenStudy (anonymous):

Just add the like term..

OpenStudy (anonymous):

\[(6x + 3x + x) - 50 = 180\] Add the terms in the brackets first...

OpenStudy (jiteshmeghwal9):

combine like terms first 6x+3x-50+x=180 10x=180+50 10x=230 x=230/10=???

OpenStudy (anonymous):

How much 6 + 3 + 1 = ???

OpenStudy (anonymous):

@jiteshmeghwal9 don't you think that you are giving the solution easily to the asker??

OpenStudy (jiteshmeghwal9):

sorry!